OpenStudy (anonymous):

rectangular prism with a square base has a lateral area of 192 square yards and a surface area of 264 square yards. What is the length of a base edge?

OpenStudy (anonymous):

Let x = the length of a side of the base, and let y = the height of the prism. LA = 4xy = 192 SA = 2x^2 + 4xy = 264 2x^2 = 72 x^2 = 36 x = 6 The length of a base edge is 6 yd.